# Representing in Low Rank I: conjugacy, topological and homological aspects

### v1

@@ 5.2. Virtually free quotients of congruence subgroups
Hence, the subgroup $\operatorname{PCon}_{n}$ is normalised by a reflection. This shows, by [Lubot96, Corollary 3.6], that $\operatorname{PCon}_{n}$ maps to a virtually free group. ∎
@@
From the expression of $\iota$, it is not hard to see that $\iota(\ker\varphi)\leq\ker\psi$. From the assumptions on $q$, i.e. that $u$ is not a square modulo $q$, it follows that $q\mathcal{I}_{u}$ is a prime ideal. This is well-known, for example see [Samuel2, § 5.4 Proposition 1]. As such, from [GurLor, Lemma 9] it follows that every torsion element of the kernel has order some power of $q$. On the other hand, by a result of Schur which is also stated in [GurLor, Theorem 14], there exists a natural number $s$ such that the orders of the torsion elements of $\operatorname{GL}_{4}(\mathcal{I}_{u})$ divide $s$.
@@ 5.3. Congruence kernel of higher modular groups
The proof proceeds as for [Lubot82, Theorem B]. Namely, reading between the lines in loc. cit., one obtains a reduction to the following claim.
@@
It follows from [Lubot82, Theorem 2.1.(a)] that the non-trivial normal subgroup $Z\cap C(\Gamma)$ of $Z\cong\widehat{F_{m}}$ contains $\widehat{F_{\omega}}$. In particular also $C(\Gamma)$ contains $\widehat{F_{\omega}}$, as desired. Thus it indeed remains to prove the claim.
@@
If $\left(\frac{u,v}{{\mathbb{Q}}_{p_{i}}}\right)\cong\operatorname{M}_{2}({\mathbb{Q}}_{p_{i}})$ is split, then $\mathcal{O}/P_{i}\cong\operatorname{M}_{2}({\mathbb{F}}_{p})$ by [Reiner, Theorem (22.4)]. Furthermore, if $\left(\frac{u,v}{{\mathbb{Q}}_{p_{i}}}\right)$ is still a division algebra, then $\mathcal{O}/P_{i}\cong{\mathbb{F}}_{p^{2}}$. In conclusion, $\operatorname{SL}_{n}(\mathcal{O}/P_{i})$ is of the form $\operatorname{SL}_{dn}({\mathbb{F}}_{p^{t}})$ for some $d,t\in{\mathbb{N}}$. It follows from [Lubot82, Lemma 2.4.(b)] that there is a finite non-abelian simple group which is not the quotient of subgroup of $\operatorname{SL}_{dn}({\mathbb{F}}_{p^{t}})$ for any $p$, yielding the desired contradiction and as such finishing the proof. ∎
@@ 6.1. Characterisation of $\operatorname{(M_{exc})}$ via largeness of (higher) modular groups
A main ingredient for the equivalence between (1) and (2) is Lemma˜2.1. In particular, we need to show that the class of virtually-$\prod\mathcal{G}_{\operatorname{vQL}}$ groups satisfies the necessary properties. The proof of [FreebyFree, Theorem 2.1] is particularly instructive for the direct factor property.
@@
Indeed, if $\operatorname{M}_{n}(D)$ is not exceptional, then $\Gamma$ satisfies property (FAb) by Margulis’s normal subgroup theorem [MargulisBook] and hence cannot satisfy largeness. If $\operatorname{M}_{n}(D)=\operatorname{M}_{2}({\mathbb{Q}}(\sqrt{-d}))$ for $d\in{\mathbb{N}}$, then it has $\operatorname{vQL}$ by [GS] (or [Lubot96, Theorem 3.7]). If $\operatorname{M}_{n}(D)=\operatorname{M}_{2}(\left(\frac{-a,-b}{{\mathbb{Q}}}\right))$, then $\operatorname{vQL}$ is given by Theorem˜5.1 and the fact that $\operatorname{vQL}$ is preserved by commensurability. This proves (6.1) and hence the moreover part.
Now we focus on the equivalence between (1) and (2). By Lemma˜6.2 the class virtually-$\prod\mathcal{G}_{\operatorname{vQL}}$ satisfies the necessary properties to apply Lemma˜2.1. Consequently, ${\mathcal{U}}(RG)$ is virtually-$\prod\mathcal{G}_{\operatorname{vQL}}$ if and only if $\operatorname{SL}_{1}(RGe)$ is either finite or virtually-$\mathcal{G}_{\operatorname{vQL}}$ for each $e\in\operatorname{PCI}(FG)$. The former only occurs if $FGe$ is a field or totally definite quaternion algebra [EricAngel1, Proposition 5.5.6]. To understand when $\operatorname{SL}_{1}(RGe)$ is virtually-$\mathcal{G}_{\operatorname{vQL}}$ one needs to distinguish whether $FGe$ is a division algebra or not.
If $FGe$ is not a division algebra, then $\operatorname{SL}_{1}(RGe)$ has $\operatorname{vQL}$ if and only if $FGe$ is an exceptional matrix algebra by (6.1). Moreover $\operatorname{SL}_{1}(RGe)$ is virtually indecomposable by [KleRio, Theorem 1]. This finishes the proof that (1) implies that $FG$ has $\operatorname{(M_{exc})}$. Therefore, now suppose that $FGe$ has is one of the possible non-commutative division algebra component of a group algebra with $\operatorname{(M_{exc})}$. Then $D$ is either exceptional of the form $\left(\frac{\zeta_{2^{t}},-3}{{\mathbb{Q}}(\zeta_{2^{t}})}\right)$ with $t\geq 4$ or it is a totally definite quaternion algebra. In the latter case ${\mathcal{U}}(\mathcal{O})$ is finite by [Kleinert]. In the former case the $S$-rank, with $S$ the set of infinite places, is at least $2$ and hence by Margulis’s normal subgroup theorem $\operatorname{SL}_{1}(\mathcal{O})$ has property (FAb) and thus not $\operatorname{vQL}$. This both finishes the proof that (1) implies (2) and the converse. ∎
@@ 6.2. Congruence kernel of group rings
The congruence kernel of an integral group ring has been investigated in [CdR1, CdR2, CdRZ]. In loc. cit. several families of finite groups $G$ are described with the property that if $C_{G}$ is infinite, then $G$ is a member of one these families. Under the assumption that ${\mathbb{Q}}G$ has no exceptional division components, an upper-bound on the virtual cohomological dimension of $C_{G}$ is obtained in [CdRZ, Theorem 5.1]. We contribute towards bounding from below, by showing that whenever ${\mathbb{Q}}G$ has an exceptional matrix component, then the congruence kernel contains a profinite free group of countable rank. This is a direct consequence of [Lubot82, Theorem B], Theorem˜5.7 and the well-behavedness of the congruence kernel with respect to direct products.
@@ 7. The blockwise Zassenhaus and subgroup isomorphism property
Three conjectures concerning finite subgroups $H$ of the normalised unit group $V({\mathbb{Z}}G)$ of an integral group ring were attributed to Zassenhaus [Zass, SehAntwerp]. The strongest (referred to as the third) Zassenhaus conjecture predicts that $H$ is conjugated over ${\mathbb{Q}}G$ to a subgroup of $G$. The second asserts the same, but only under the hypothesis that $|H|=|G|$, and the first (and weakest) concerns the case that $H$ is cyclic. All of these conjectures have been disproven [Scott, Rogg, EisMar], transforming them into the question for which $H$ and $G$ the above conjugation property holds. In Section˜7.1 we introduce a blockwise variant of the conjectures, i.e. investigating whether the image of $H$ under any irreducible $\mathbb{Q}$}-representation $\rho$ of $G$ is conjugated in $\rho({\mathbb{Q}}G)$ to a subgroup of $\rho(G)$. Furthermore, we define the Zassenhaus property for semisimple $F$-algebras with $F$ a field of characteristic 0. The rest of the section is devoted to the investigation of the Zassenhaus property for exceptional matrix algebras. Among other things, in Section˜7.4 we describe the conjugacy classes of finite subgroups of exceptional matrix algebras.
@@ 7.1. The Zassenhaus property for semisimple algebras
The subgroup isomorphism property for the sets $\mathcal{G}_{i}$, with $i=1,2,3$, has also been investigated, see [Margolis, PdRV] and references therein. The subgroup isomorphism property for the set $\mathcal{G}_{2}$ and twisted group rings has recently also received considerable attention, see [MaSc, MaSc2, MaSc3, HKSa, HKSi].
Also, implicitly the blockwise Zassenhaus property has been considered in [HerSin, Lemma 3].
@@
Suppose first that $H\leqslant\Gamma$. Denote $H=\langle h\rangle$. Then from [RoggenkampTaylor, Theorem V.1] it follows that there is a bijection $\varphi\colon\Gamma\to G$ such that for all $x\in\Gamma$ the following holds: (i) $var\phi(x)$ and $x$ have the same order, (ii) $\varphi(x)^{n}$ and $\varphi(x^{n})$ are conjugate for all $n\in{\mathbb{Z}}$, and (iii) $\chi(x)=\chi(\phi(x))$ for any irreducible $K$-character, with $K$ some field containing $F$. Then fixing $x\in\Gamma$ such that $\phi(x)=h$, the restriction $\varphi|_{\langle x\rangle}\colon\langle x\rangle\to H$ is a group isomorphism. Moreover, for any $\chi$ as above and any $j\in{\mathbb{Z}}$, $\chi(x^{j})=\chi(h^{j})$. It follows from [^16] Note that the statement is only formulated for $F={\mathbb{Q}}$, but the proof can also be generalised to this context, by replacing $\mathbb{C}$} in the proof of [SehgalBook93, Lemma 37.6] with a splitting field of $G$ containing $F$. [SehgalBook93, Lemma 37.6] that $x$ and $h$ are conjugate over $FG$.
@@
One could wonder whether there is a version of Proposition˜7.5 with $\Gamma^{y}$ replaced by any group basis of ${\mathbb{Q}}G$. This however seems not to be possible. Concretely, Leo Margolis communicated to the authors that the cyclic subgroup $C_{n}$ serving as counterexample to the first Zassenhaus conjecture (see [EisMar]) is expected not to be contained in any group basis of ${\mathbb{Q}}G$. It would be interesting to investigate whether there is some $e\in\operatorname{PCI}({\mathbb{Q}}G)$ such that $C_{n}e$ is contained in no spanning subgroup of ${\mathcal{U}}({\mathbb{Q}}Ge)$ of the form $\Gamma^{y}$ with $\Gamma\leq V({\mathbb{Z}}G)$ and $y\in{\mathbb{Q}}Ge$. In particular, the following is natural to ask.
@@
Does the counterexample to the first Zassenhaus conjecture in [EisMar] satisfy the blockwise Zassenhaus property relative to cyclic subgroups?
@@
The spectrum problem is the weaker version of the first Zassenhaus conjecture where ‘conjugate’ is replaced by ‘isomorphic to’. In contrast to the Zassenhaus conjecture, the spectrum problem is still open. In [HertOrder], it was notably proven to be true for $G$ solvable. Moreover, if $R$ is the ring of integers in a number field, a theorem of Cohn–Livingstone [CL] yields that the unit group $V(RG)$ contains an element of prime power order $p^{n}$ if and only if $G$ does. In particular,
@@ 7.2. Somme illustrative examples
with $C_{2}\times C_{2}=\langle a^{o(a)/2},b\rangle$, for $a$ and $b$ such that $D_{n}=\langle a,b\mid a^{n/2}=b^{2}=1,bab=a^{-1}\rangle$. By Bass–Serre theory (see e.g. [Serre, I.4, Theorem 8]), under the aforementioned isomorphism, an arbitrary finite subgroup $H$ of $\operatorname{GL}_{2}({\mathbb{Z}})$ is conjugate to a subgroup of $D_{8}$ or of $D_{12}$. In particular, if $H$ is not a $2$-group and $|H|\neq 12$, then $H$ is conjugated to a subgroup of $D_{6}$.
@@
We would like to emphasise that in this example one has an especially nice Zassenhaus property since the conjugation is already over the order. However, this is specific to $\operatorname{M}_{2}({\mathbb{Q}})$, as the upcoming example already illustrates well, see Remark˜7.13. Furthermore, Proposition˜7.11 was already implicit in [J1, Theorem 1.5] were it is proven (using other terminology) that if all the non-division components of ${\mathbb{Q}}G$ are isomorphic to $\operatorname{M}_{2}({\mathbb{Q}})$, then ${\mathbb{Z}}G$ enjoys the blockwise Zassenhaus property relative to $G$ and the set of all finite subgroups. Moreover, conjugation can be realised over the maximal order. Importantly, the latter is really a property of the projections. Indeed, it is classical that ${\mathbb{Z}}S_{3}$ satisfies the third Zassenhaus conjecture, however there is a finite subgroup of $V({\mathbb{Z}}S_{3})$ which is conjugated to a subgroup of $S_{3}$ over ${\mathbb{Q}}S_{3}$, but not over ${\mathbb{Z}}S_{3}$.
@@
In [Hat], Hatcher gave a decomposition as an amalgamated product for $\operatorname{PGL}_{2}(\mathcal{I}_{2})$, where $\mathcal{I}_{2}$ is the ring of integers of ${\mathbb{Q}}(\sqrt{-2})$. This decomposition can be adapted to yield the following for $\operatorname{GL}_{2}(\mathcal{I}_{2})$, see Remark˜7.37 for details:
@@
Next suppose that ${\mathbb{Q}}Ge$ is a quaternion algebra. Then $Ge$ is one of the groups mentioned in Lemma˜7.35. Now recall that the third Zassenhaus conjecture was proven for $Q_{4m}$ in [CMdR], for $\operatorname{SL}_{2}({\mathbb{F}}_{5})$ in [DJPm, Theorem 4.3] and in [DJ, Theorem 4.7] for $\operatorname{SU}_{2}({\mathbb{F}}_{3})$. In these cases, using (7.7) and (7.8), the desired conjugation is inherited from $V({\mathbb{Z}}[Ge])$.
It remains to consider the case that $Ge\cong\operatorname{SL}_{2}({\mathbb{F}}_{3})$. In this case the first Zassenhaus conjecture was obtained in [HK], and $\operatorname{Span}_{{\mathbb{Q}}}\{Ge\}=\left(\frac{-1,-1}{{\mathbb{Q}}}\right)$. More precisely, $Ge$ is isomorphic to the unit group of the Hurwitz quaternions, which is up to conjugation the unique maximal order in $\left(\frac{-1,-1}{{\mathbb{Q}}}\right)$ and thus $He$ is conjugated to a subgroup of $Ge$, finishing the proof. ∎
@@
This example will require combining the methods from the previous examples. For $\operatorname{GL}_{2}({\mathbb{Z}}[\sqrt{-1}])$, it is not possible to follow the exact same strategy as for $\operatorname{GL}_{2}({\mathbb{Z}})$ and $\operatorname{GL}_{2}(\mathcal{I}_{2})$, since $\operatorname{GL}_{2}({\mathbb{Z}}[\sqrt{-1}])$ has Serre’s property (FA) [BJJKT, Theorem 5.1] and thus cannot be decomposed through amalgamation. Nevertheless, in [FineBook, Theorem 4.4.1] an amalgamated decomposition was constructed for $\operatorname{PSL}_{2}({\mathbb{Z}}[\sqrt{-1}])$ from which one readily deduces the following decomposition for $\operatorname{SL}_{2}({\mathbb{Z}}[\sqrt{-1}])$:
@@
with $Q$ a subgroup of $\operatorname{PGL}_{2}({\mathbb{Z}}[\sqrt{-1}])$. Subgroups of the latter are well-known (e.g. see [Beau]): $C_{n}$ with $n\mid 4,6$, $D_{6}$, $D_{8}$, $A_{4}$ and $S_{4}$. A computation of their second group cohomology group with coefficients in $C_{2}$ or $C_{4}$ would yield the same conclusion.
@@
By the linear independence of finite subgroups of $V({\mathbb{Z}}G)$, it follows that $|H|$ divides $|C_{4}\times S_{3}|=24$. If $|H|=24,$ then $H\cong C_{4}\times S_{3}$ as $C_{4}\times S_{3}$ is metabelian and hence satisfies the isomorphism problem by Whitcomb’s theorem. Next let $|H|\mid 8$ or $12$. Moreover, since $C_{4}\times S_{3}$ is solvable, it follows from [HertOrder] that $\operatorname{spec}(H)\subseteq\operatorname{spec}(C_{4}\times S_{3})=\{n\,:\,n\mid 12\}$. This in particular solves the case that $H$ is cyclic. Next recall that the subgroup isomorphism problem is known for subgroups of the form $C_{2}\times C_{2}$ ([HertOrder2]), $C_{2}\times C_{4}$ ([MarSIP]) and $C_{p}\times C_{q}^{\ell}$ for $p\neq q$ primes ([HertOrder]). Hence for $H$ abelian, the only remaining case to investigate is $H\cong C_{2}^{3}$. However, this is not possible since $H$ embeds into ${\mathbb{Z}}Ge$, where the maximal rank of an elementary abelian $2$-subgroup is $2$.

### v2

@@ 5.2. Virtually free quotients of congruence subgroups
Hence, the subgroup $\operatorname{PCon}_{n}$ is normalised by a reflection. This shows, by [59, Corollary 3.6], that $\operatorname{PCon}_{n}$ maps to a virtually free group. ∎
@@
From the expression of $\iota$, it is not hard to see that $\iota(\ker\varphi)\leq\ker\psi$. From the assumptions on $q$, i.e. that $u$ is not a square modulo $q$, it follows that $q\mathcal{I}_{u}$ is a prime ideal. This is well-known, for example see [76, § 5.4 Proposition 1]. As such, from [33, Lemma 9] it follows that every torsion element of the kernel has order some power of $q$. On the other hand, by a result of Schur which is also stated in [33, Theorem 14], there exists a natural number $s$ such that the orders of the torsion elements of $\operatorname{GL}_{4}(\mathcal{I}_{u})$ divide $s$.
@@ 5.3. Congruence kernel of higher modular groups
The proof proceeds as for [58, Theorem B]. Namely, reading between the lines in loc. cit., one obtains a reduction to the following claim.
@@
It follows from [58, Theorem 2.1.(a)] that the non-trivial normal subgroup $Z\cap C(\Gamma)$ of $Z\cong\widehat{F_{m}}$ contains $\widehat{F_{\omega}}$. In particular also $C(\Gamma)$ contains $\widehat{F_{\omega}}$, as desired. Thus it indeed remains to prove the claim.
@@
If $\left(\frac{u,v}{{\mathbb{Q}}_{p_{i}}}\right)\cong\operatorname{M}_{2}({\mathbb{Q}}_{p_{i}})$ is split, then $\mathcal{O}/P_{i}\cong\operatorname{M}_{2}({\mathbb{F}}_{p})$ by [73, Theorem (22.4)]. Furthermore, if $\left(\frac{u,v}{{\mathbb{Q}}_{p_{i}}}\right)$ is still a division algebra, then $\mathcal{O}/P_{i}\cong{\mathbb{F}}_{p^{2}}$. In conclusion, $\operatorname{SL}_{n}(\mathcal{O}/P_{i})$ is of the form $\operatorname{SL}_{dn}({\mathbb{F}}_{p^{t}})$ for some $d,t\in{\mathbb{N}}$. It follows from [58, Lemma 2.4.(b)] that there is a finite non-abelian simple group which is not the quotient of subgroup of $\operatorname{SL}_{dn}({\mathbb{F}}_{p^{t}})$ for any $p$, yielding the desired contradiction and as such finishing the proof. ∎
@@ 6.1. Characterisation of $\operatorname{(M_{exc})}$ via largeness of (higher) modular groups
A main ingredient for the equivalence between (1) and (2) is Lemma˜2.1. In particular, we need to show that the class of virtually-$\prod\mathcal{G}_{\operatorname{vQL}}$ groups satisfies the necessary properties. The proof of [50, Theorem 2.1] is particularly instructive for the direct factor property.
@@
Indeed, if $\operatorname{M}_{n}(D)$ is not exceptional, then $\Gamma$ satisfies property (FAb) by Margulis’s normal subgroup theorem [66] and hence cannot satisfy largeness. If $\operatorname{M}_{n}(D)=\operatorname{M}_{2}({\mathbb{Q}}(\sqrt{-d}))$ for $d\in{\mathbb{N}}$, then it has $\operatorname{vQL}$ by [32] (or [59, Theorem 3.7]). If $\operatorname{M}_{n}(D)=\operatorname{M}_{2}(\left(\frac{-a,-b}{{\mathbb{Q}}}\right))$, then $\operatorname{vQL}$ is given by Theorem˜5.1 and the fact that $\operatorname{vQL}$ is preserved by commensurability. This proves (6.1) and hence the moreover part.
Now we focus on the equivalence between (1) and (2). By Lemma˜6.2 the class virtually-$\prod\mathcal{G}_{\operatorname{vQL}}$ satisfies the necessary properties to apply Lemma˜2.1. Consequently, ${\mathcal{U}}(RG)$ is virtually-$\prod\mathcal{G}_{\operatorname{vQL}}$ if and only if $\operatorname{SL}_{1}(RGe)$ is either finite or virtually-$\mathcal{G}_{\operatorname{vQL}}$ for each $e\in\operatorname{PCI}(FG)$. The former only occurs if $FGe$ is a field or totally definite quaternion algebra [47, Proposition 5.5.6]. To understand when $\operatorname{SL}_{1}(RGe)$ is virtually-$\mathcal{G}_{\operatorname{vQL}}$ one needs to distinguish whether $FGe$ is a division algebra or not.
If $FGe$ is not a division algebra, then $\operatorname{SL}_{1}(RGe)$ has $\operatorname{vQL}$ if and only if $FGe$ is an exceptional matrix algebra by (6.1). Moreover $\operatorname{SL}_{1}(RGe)$ is virtually indecomposable by [55, Theorem 1]. This finishes the proof that (1) implies that $FG$ has $\operatorname{(M_{exc})}$. Therefore, now suppose that $FGe$ has is one of the possible non-commutative division algebra component of a group algebra with $\operatorname{(M_{exc})}$. Then $D$ is either exceptional of the form $\left(\frac{\zeta_{2^{t}},-3}{{\mathbb{Q}}(\zeta_{2^{t}})}\right)$ with $t\geq 4$ or it is a totally definite quaternion algebra. In the latter case ${\mathcal{U}}(\mathcal{O})$ is finite by [54]. In the former case the $S$-rank, with $S$ the set of infinite places, is at least $2$ and hence by Margulis’s normal subgroup theorem $\operatorname{SL}_{1}(\mathcal{O})$ has property (FAb) and thus not $\operatorname{vQL}$. This both finishes the proof that (1) implies (2) and the converse. ∎
@@ 6.2. Congruence kernel of group rings
The congruence kernel of an integral group ring has been investigated in [9, 10, 13]. In loc. cit. several families of finite groups $G$ are described with the property that if $C_{G}$ is infinite, then $G$ is a member of one these families. Under the assumption that ${\mathbb{Q}}G$ has no exceptional division components, an upper-bound on the virtual cohomological dimension of $C_{G}$ is obtained in [13, Theorem 5.1]. We contribute towards bounding from below, by showing that whenever ${\mathbb{Q}}G$ has an exceptional matrix component, then the congruence kernel contains a profinite free group of countable rank. This is a direct consequence of [58, Theorem B], Theorem˜5.7 and the well-behavedness of the congruence kernel with respect to direct products.
@@ 7. The blockwise Zassenhaus and subgroup isomorphism property
Three conjectures concerning finite subgroups $H$ of the normalised unit group $V({\mathbb{Z}}G)$ of an integral group ring were attributed to Zassenhaus [88, 80]. The strongest (referred to as the third) Zassenhaus conjecture predicts that $H$ is conjugated over ${\mathbb{Q}}G$ to a subgroup of $G$. The second asserts the same, but only under the hypothesis that $|H|=|G|$, and the first (and weakest) concerns the case that $H$ is cyclic. All of these conjectures have been disproven [77, 74, 24], transforming them into the question for which $H$ and $G$ the above conjugation property holds. In Section˜7.1 we introduce a blockwise variant of the conjectures, i.e. investigating whether the image of $H$ under any irreducible $\mathbb{Q}$}-representation $\rho$ of $G$ is conjugated in $\rho({\mathbb{Q}}G)$ to a subgroup of $\rho(G)$. Furthermore, we define the Zassenhaus property for semisimple $F$-algebras with $F$ a field of characteristic 0. The rest of the section is devoted to the investigation of the Zassenhaus property for exceptional matrix algebras. Among other things, in Section˜7.4 we describe the conjugacy classes of finite subgroups of exceptional matrix algebras.
@@ 7.1. The Zassenhaus property for semisimple algebras
The subgroup isomorphism property for the sets $\mathcal{G}_{i}$, with $i=1,2,3$, has also been investigated, see [60, 68] and references therein. The subgroup isomorphism property for the set $\mathcal{G}_{2}$ and twisted group rings has recently also received considerable attention, see [63, 64, 65, 35, 51].
Also, implicitly the blockwise Zassenhaus property has been considered in [36, Lemma 3].
@@
Suppose first that $H\leqslant\Gamma$. Denote $H=\langle h\rangle$. Then from [75, Theorem V.1] it follows that there is a bijection $\varphi\colon\Gamma\to G$ such that for all $x\in\Gamma$ the following holds: (i) $var\phi(x)$ and $x$ have the same order, (ii) $\varphi(x)^{n}$ and $\varphi(x^{n})$ are conjugate for all $n\in{\mathbb{Z}}$, and (iii) $\chi(x)=\chi(\phi(x))$ for any irreducible $K$-character, with $K$ some field containing $F$. Then fixing $x\in\Gamma$ such that $\phi(x)=h$, the restriction $\varphi|_{\langle x\rangle}\colon\langle x\rangle\to H$ is a group isomorphism. Moreover, for any $\chi$ as above and any $j\in{\mathbb{Z}}$, $\chi(x^{j})=\chi(h^{j})$. It follows from [^16] Note that the statement is only formulated for $F={\mathbb{Q}}$, but the proof can also be generalised to this context, by replacing $\mathbb{C}$} in the proof of [78, Lemma 37.6] with a splitting field of $G$ containing $F$. [78, Lemma 37.6] that $x$ and $h$ are conjugate over $FG$.
@@
One could wonder whether there is a version of Proposition˜7.5 with $\Gamma^{y}$ replaced by any group basis of ${\mathbb{Q}}G$. This however seems not to be possible. Concretely, Leo Margolis communicated to the authors that the cyclic subgroup $C_{n}$ serving as counterexample to the first Zassenhaus conjecture (see [24]) is expected not to be contained in any group basis of ${\mathbb{Q}}G$. It would be interesting to investigate whether there is some $e\in\operatorname{PCI}({\mathbb{Q}}G)$ such that $C_{n}e$ is contained in no spanning subgroup of ${\mathcal{U}}({\mathbb{Q}}Ge)$ of the form $\Gamma^{y}$ with $\Gamma\leq V({\mathbb{Z}}G)$ and $y\in{\mathbb{Q}}Ge$. In particular, the following is natural to ask.
@@
Does the counterexample to the first Zassenhaus conjecture in [24] satisfy the blockwise Zassenhaus property relative to cyclic subgroups?
@@
The spectrum problem is the weaker version of the first Zassenhaus conjecture where ‘conjugate’ is replaced by ‘isomorphic to’. In contrast to the Zassenhaus conjecture, the spectrum problem is still open. In [38], it was notably proven to be true for $G$ solvable. Moreover, if $R$ is the ring of integers in a number field, a theorem of Cohn–Livingstone [14] yields that the unit group $V(RG)$ contains an element of prime power order $p^{n}$ if and only if $G$ does. In particular,
@@ 7.2. Somme illustrative examples
with $C_{2}\times C_{2}=\langle a^{o(a)/2},b\rangle$, for $a$ and $b$ such that $D_{n}=\langle a,b\mid a^{n/2}=b^{2}=1,bab=a^{-1}\rangle$. By Bass–Serre theory (see e.g. [82, I.4, Theorem 8]), under the aforementioned isomorphism, an arbitrary finite subgroup $H$ of $\operatorname{GL}_{2}({\mathbb{Z}})$ is conjugate to a subgroup of $D_{8}$ or of $D_{12}$. In particular, if $H$ is not a $2$-group and $|H|\neq 12$, then $H$ is conjugated to a subgroup of $D_{6}$.
@@
We would like to emphasise that in this example one has an especially nice Zassenhaus property since the conjugation is already over the order. However, this is specific to $\operatorname{M}_{2}({\mathbb{Q}})$, as the upcoming example already illustrates well, see Remark˜7.13. Furthermore, Proposition˜7.11 was already implicit in [42, Theorem 1.5] were it is proven (using other terminology) that if all the non-division components of ${\mathbb{Q}}G$ are isomorphic to $\operatorname{M}_{2}({\mathbb{Q}})$, then ${\mathbb{Z}}G$ enjoys the blockwise Zassenhaus property relative to $G$ and the set of all finite subgroups. Moreover, conjugation can be realised over the maximal order. Importantly, the latter is really a property of the projections. Indeed, it is classical that ${\mathbb{Z}}S_{3}$ satisfies the third Zassenhaus conjecture, however there is a finite subgroup of $V({\mathbb{Z}}S_{3})$ which is conjugated to a subgroup of $S_{3}$ over ${\mathbb{Q}}S_{3}$, but not over ${\mathbb{Z}}S_{3}$.
@@
In [34], Hatcher gave a decomposition as an amalgamated product for $\operatorname{PGL}_{2}(\mathcal{I}_{2})$, where $\mathcal{I}_{2}$ is the ring of integers of ${\mathbb{Q}}(\sqrt{-2})$. This decomposition can be adapted to yield the following for $\operatorname{GL}_{2}(\mathcal{I}_{2})$, see Remark˜7.37 for details:
@@
Next suppose that ${\mathbb{Q}}Ge$ is a quaternion algebra. Then $Ge$ is one of the groups mentioned in Lemma˜7.35. Now recall that the third Zassenhaus conjecture was proven for $Q_{4m}$ in [11], for $\operatorname{SL}_{2}({\mathbb{F}}_{5})$ in [22, Theorem 4.3] and in [21, Theorem 4.7] for $\operatorname{SU}_{2}({\mathbb{F}}_{3})$. In these cases, using (7.7) and (7.8), the desired conjugation is inherited from $V({\mathbb{Z}}[Ge])$.
It remains to consider the case that $Ge\cong\operatorname{SL}_{2}({\mathbb{F}}_{3})$. In this case the first Zassenhaus conjecture was obtained in [41], and $\operatorname{Span}_{{\mathbb{Q}}}\{Ge\}=\left(\frac{-1,-1}{{\mathbb{Q}}}\right)$. More precisely, $Ge$ is isomorphic to the unit group of the Hurwitz quaternions, which is up to conjugation the unique maximal order in $\left(\frac{-1,-1}{{\mathbb{Q}}}\right)$ and thus $He$ is conjugated to a subgroup of $Ge$, finishing the proof. ∎
@@
This example will require combining the methods from the previous examples. For $\operatorname{GL}_{2}({\mathbb{Z}}[\sqrt{-1}])$, it is not possible to follow the exact same strategy as for $\operatorname{GL}_{2}({\mathbb{Z}})$ and $\operatorname{GL}_{2}(\mathcal{I}_{2})$, since $\operatorname{GL}_{2}({\mathbb{Z}}[\sqrt{-1}])$ has Serre’s property (FA) [4, Theorem 5.1] and thus cannot be decomposed through amalgamation. Nevertheless, in [27, Theorem 4.4.1] an amalgamated decomposition was constructed for $\operatorname{PSL}_{2}({\mathbb{Z}}[\sqrt{-1}])$ from which one readily deduces the following decomposition for $\operatorname{SL}_{2}({\mathbb{Z}}[\sqrt{-1}])$:
@@
with $Q$ a subgroup of $\operatorname{PGL}_{2}({\mathbb{Z}}[\sqrt{-1}])$. Subgroups of the latter are well-known (e.g. see [7]): $C_{n}$ with $n\mid 4,6$, $D_{6}$, $D_{8}$, $A_{4}$ and $S_{4}$. A computation of their second group cohomology group with coefficients in $C_{2}$ or $C_{4}$ would yield the same conclusion.
@@
By the linear independence of finite subgroups of $V({\mathbb{Z}}G)$, it follows that $|H|$ divides $|C_{4}\times S_{3}|=24$. If $|H|=24,$ then $H\cong C_{4}\times S_{3}$ as $C_{4}\times S_{3}$ is metabelian and hence satisfies the isomorphism problem by Whitcomb’s theorem. Next let $|H|\mid 8$ or $12$. Moreover, since $C_{4}\times S_{3}$ is solvable, it follows from [38] that $\operatorname{spec}(H)\subseteq\operatorname{spec}(C_{4}\times S_{3})=\{n\,:\,n\mid 12\}$. This in particular solves the case that $H$ is cyclic. Next recall that the subgroup isomorphism problem is known for subgroups of the form $C_{2}\times C_{2}$ ([40]), $C_{2}\times C_{4}$ ([61]) and $C_{p}\times C_{q}^{\ell}$ for $p\neq q$ primes ([38]). Hence for $H$ abelian, the only remaining case to investigate is $H\cong C_{2}^{3}$. However, this is not possible since $H$ embeds into ${\mathbb{Z}}Ge$, where the maximal rank of an elementary abelian $2$-subgroup is $2$.